# HybridNeRF: Efficient Neural Rendering via Adaptive Volumetric Surfaces

### v1

@@ Abstract
Neural radiance fields provide state-of-the-art view synthesis quality but tend to be slow to render. One reason is that they make use of volume rendering, thus requiring many samples (and model queries) per ray at render time. Although this representation is flexible and easy to optimize, most real-world objects can be modeled more efficiently with surfaces instead of volumes, requiring far fewer samples per ray. This observation has spurred considerable progress in surface representations such as signed distance functions, but these may struggle to model semi-opaque and thin structures. We propose a method, HybridNeRF, that leverages the strengths of both representations by rendering most objects as surfaces while modeling the (typically) small fraction of challenging regions volumetrically. We evaluate HybridNeRF against the challenging Eyeful Tower dataset [35] along with other commonly used view synthesis datasets. When comparing to state-of-the-art baselines, including recent rasterization-based approaches, we improve error rates by 15–30% while achieving real-time framerates (at least 36 FPS) for virtual-reality resolutions (2K×$\times$2K).
@@ 1 Introduction
Figure 2: Approach. In the first phase of our pipeline (a), we train a VolSDF-like [36] model with distance-adjusted Eikonal loss to model backgrounds without a separate NeRF (Section 3.3). We then crucially transition from a uniform surfaceness parameter $\beta$ to position-dependent $\beta(\boldsymbol{x})$ values to model most of the scene as thin surfaces (needing few samples) without degrading quality near fine and semi-opaque structures (b). Since our model behaves as a valid SDF in $>$95% of the scene, we use sphere tracing at render time (c) along with lower-level optimizations (hardware texture interpolation) to query each sample as efficiently as possible.
@@
While surface-based neural fields are convenient for rendering, they often struggle to reconstruct scenes with thin structures or view-dependent effects, such as reflections and translucency. This is one reason that surfaces are often transformed into volumetric models for rendering [36]. A crucial transformation parameter is a scalar temperature $\beta$ that is used to convert a $\beta$-scaled signed distance value into a density. Higher temperatures tend to produce an ideal binary occupancy field that can improve rendering speed but can struggle for challenging regions as explained above. Lower temperatures allow the final occupancy field to remain flexible, whereby the $\beta$-scaled SDF essentially acts as a reparameterization of the underlying occupancy field. As such, we refer to $\beta$ as the surfaceness of the underlying scene (see Figure 1). Prior work treats $\beta$ as a global parameter that is explicitly scheduled or learned via gradient descent [36]. We learn it in a spatially adaptive manner.
@@ 2 Related Work
Some of the earliest NeRF acceleration methods store precomputed non-view dependent model outputs, such as spherical harmonics coefficients, into finite-resolution structures [40, 11, 7, 5]. These outputs are combined with viewing direction to compute the final radiance at render time, bypassing the original model entirely. Although these methods render extremely quickly (some $>$200 FPS [7]), they are limited by the finite capacity of the caching structure and cannot capture fine details at room scale.
@@
Several methods [23, 33, 36] derive density values from the outputs of a signed distance function, which are then rendered volumetrically as in NeRF. These hybrid representations retain NeRF’s ease of optimization while learning better surface geometry. Follow-up works [37, 8] bake the resulting surface geometry into a mesh that is further optimized and simplified. Similar to early voxel-baking approaches, these methods render quickly ($>$70 FPS) but are limited by the capacity of the mesh and texture, and thus struggle to model thin structures, transparency, and view-dependent effects. We train a similar SDF representation in our method but continue to use the base neural model at render time.
@@ 3 Method
Given a collection of RGB images and camera poses, our goal is to learn a 3D representation that generates novel views at VR resolution (at least 2K×$\times$2K pixels) in real-time (at least 36 FPS), while achieving a high degree of visual fidelity. As we target captures taken under real-world conditions, our representation must be able to account for inconsistencies across training images due to lighting changes and shadows (even in “static” scenes). Our approach builds upon NeRF’s raycasting paradigm, which can generate highly photorealistic renderings, and improves upon its efficiency. Since most of the world consists of surfaces, we develop a representation that can render surfaces with few samples and without degrading the rest of the scene. We outline our approach in Figure 2 and present our model architecture and the first training stage in Section 3.1, which is followed by finetuning of our model to accelerate rendering without compromising quality in Section 3.2. We discuss how to model unbounded scenes in Section 3.3 and present final render-time optimizations in Section 3.4.
@@ 3.1 Representation
NeRF [20] represents a scene as a continuous volumetric radiance field that encodes the scene’s geometry and view-dependent appearance within the weights of a model. NeRF renders pixels by sampling positions $\boldsymbol{x}_{i}$ along the corresponding camera ray, querying the MLP to obtain density and color values, $\sigma_{i}\coloneqq\sigma(\boldsymbol{x}_{i})$ and $\mathbf{c}_{i}\coloneqq\boldsymbol{c}(\boldsymbol{x}_{i},\boldsymbol{d}_{r})$, respectively (with $\boldsymbol{d}_{r}$ as the ray direction). The density values $\sigma_{i}$ are converted into opacity values $\alpha_{i}\coloneqq 1-\exp(-\sigma_{i}\delta_{i})$, where $\delta_{i}$ is the distance between samples. The final ray color $\hat{\boldsymbol{c}}_{r}\coloneqq\sum_{i=0}^{N-1}\boldsymbol{c}_{i}w_{i}$ is obtained as a convex combination of the color samples $\boldsymbol{c}_{i}$ with weights $w_{i}\coloneqq\exp(-\!\sum_{j=0}^{i-1}\!\sigma_{j}\delta_{j})\alpha_{i}$. The training process optimizes the model by sampling batches of image pixels and minimizing the L2 reconstruction loss. We refer to Mildenhall et al. [20] for details.
@@
Figure 3: Surfaces. Since NeRF directly predicts density, it often ‘cheats’ by modeling specular surfaces such as floors as semi-transparent volumes that require many samples per ray (heatmaps shown on the right). Methods that derive density from signed distances, such as ours, improve surface geometry and appearance while using fewer samples per ray.
@@
The original NeRF representation has the flexibility of representing semi-transparent surfaces, for the density field is not forced to saturate. However, the model often abuses this property by generating semi-transparent volumes to mimic reflections and other view-dependent effects (Figure 3). This hampers our goal of minimizing the samples per ray needed for rendering.
To address this problem, surface–volume representations [23, 36, 33] learn well-defined surfaces by interpreting MLP outputs $f(\boldsymbol{x})$ as a signed distance field (SDF) such that scene surfaces are represented by the zero-level set of the function $f$.
@@
where $\beta(\boldsymbol{x})>0$ determines the surfaceness of point $\boldsymbol{x}$, i.e. how concentrated the density should be around the zero level set of $f$, and $\Psi$ is the cumulative distribution function of a standard Laplace distribution:
@@
In prior works, the surfaceness $\beta(\boldsymbol{x})$ is independent of position $\boldsymbol{x}$, and is either learned or annealed during training [37]. We instead consider a surfaceness field implemented as a $512^{3}$ grid of values that is queried via nearest-neighbor interpolation. We constrain the surfaceness parameters to be globally uniform during the first phase of training and allow them to diverge spatially during the finetuning stage (Section 3.2).
@@
We use dense multi-resolution 3D feature grids in combination with multi-resolution triplanes [3, 6] for featurizing 3D sample locations. We predict color $\boldsymbol{c}$ and signed distance $f$ using separate grids, each followed by an MLP, and use a small proposal network similar to the one used by Nerfacto [31] to improve sampling efficiency. For a given 3D point, we fetch $K\!=\!4$ features per level from (1) the 3D feature grids at 3 resolution levels (1283, 2563 and 5123) via trilinear interpolation, and (2) from triplanes of resolutions 1282 to 8,1922 via bilinear interpolation. We sum the features across levels (instead of concatenation [21, 6]), and concatenate the summed features from the 3D grid to those from the 3 triplanes to obtain a $164K\!=\!$ 16-dimensional MLP input. We encode viewing direction through spherical harmonics (up to the 4th degree) as an auxiliary input to the color MLP. As our feature grid is multi-resolution, we handle aliasing as in VR-NeRF [35] by deriving the pixel radius $p(\boldsymbol{x})$ of each sample, and dampening features whose size is less than $2p(\boldsymbol{x})$ (Appendix B).
@@ 3.2 Finetuning
The first stage of our pipeline uses a global surfaceness value $\beta(\boldsymbol{x})=\bar{\beta}$ for all $\boldsymbol{x}$, similar to existing approaches [33, 36]. As $\bar{\beta}$ increases, the density $\sigma_{\text{SDF}}$ in free-space areas converges to zero (Equation 3), thus reducing the required number of samples per ray. However, uniformly increasing this scene-wide parameter degrades the rendering quality near fine-grained and transparent structures (see Figure 4).
@@
Figure 5: Spatially adaptive surfaceness. We make $\beta(\boldsymbol{x})$ spatially adaptive by means of a $512^{3}$ voxel grid that we increase during the finetuning stage. We track Eikonal loss as we increase surfaceness as it is highest near object boundaries and semi-transparent surfaces (top-right) that degrade when surfaceness is too high (Figure 4). We stop increasing surfaceness in regions that cross a given threshold.
We overcome this limitation by making $\beta(\boldsymbol{x})$ spatially adaptive via a $512^{3}$ voxel grid. We collect per-sample triplets $(\boldsymbol{x},\eta,w)$ rendered during the finetuning process, accumulate them over multiple training iterations (5,000), and partition them across the voxels of the surfaceness grid. Let $\Lambda_{\boldsymbol{v}}$ be the subset associated with voxel $\boldsymbol{v}$ corresponding to $\beta_{\boldsymbol{v}}$. We increase $\beta_{\boldsymbol{v}}$ by a fixed increment (100) if:
@@
where $\bar{\gamma}\coloneqq 0.25$ is a predefined threshold. Figure 5 illustrates our approach.
@@
We find it important to adopt a large 256 channel-wide MLP to represent the signed distance $f$ during the initial training phase in order to preserve accurate scene geometry. However, we later distill $f$ into a much smaller 16-wide network ($f_{\text{small}})$. We do so by sampling random rays from our training set for 5,000 iterations and minimizing the absolute difference between $f(\boldsymbol{x}_{i})$ and $f_{\text{small}}(\boldsymbol{x}_{i})$ for every sampled point:
@@ 3.3 Backgrounds
Many scenes we wish to reconstruct contain complex backgrounds that surface–volume methods struggle to replicate [36, 33, 17]. BakedSDF [37] defines a contraction space [2] in which the Eikonal loss of Equation 1 is applied. However, we found this to negatively impact foreground quality. Other approaches use separate NeRF background models [42], which effectively doubles inference and memory costs, and makes them ill-suited for real-time rendering.
@@
We discuss how to make a single MLP behave as an approximate SDF in the foreground and a volumetric model in the background. Both types of NeRF derive density $\sigma$ by applying a non-linearity to the output of an MLP. Our insight is that although the original NeRF uses ReLU, any non-linear mapping to $\mathbb{R}^{+}$ may be used in practice, including our scaled CDF $\Psi$ ($\beta$ omitted without loss of generality). Since $\Psi$ is invertible (as it is a CDF), $\sigma(\boldsymbol{x})$ and $\Psi(f(\boldsymbol{x}))$ are functionally equivalent as there exists an $f$ such that $\Psi(f(\boldsymbol{x}))$ = $\sigma(\boldsymbol{x})$ for any given point $\boldsymbol{x}$. Put otherwise, it is the Eikonal regularization that causes the divergence in behavior between both methods — in its absence, an “SDF” MLP is free to behave exactly as the density MLP in the original NeRF!
@@
We use a distance-adjusted Eikonal loss during training by using per-sample loss weights $\eta_{i}=\frac{1}{d_{i}^{2}}$ (where $d_{i}$ is the metric distance along the ray of sample $\boldsymbol{x}_{i}$) instead of commonly-used uniform weights ($\eta_{i}=1$) to downweight the loss applied to far-field regions. Intuitively, this encourages our method to behave as a valid SDF in the foreground (with well-defined surfaces) and more like NeRF in the background (to enable accurate reconstruction) without the need for separate foreground and background models. Figure 6 illustrates the different approaches.
@@ 3.4 Real-Time Rendering
Volumetric methods that use occupancy grids [e.g. 21, 27] sample within occupied voxels using a given step size. This hyperparameter must be carefully tuned to strike the proper balance between quality (not skipping thin surfaces) and performance (not excessively sampling empty space). Modeling an SDF allows us to sample more efficiently by advancing toward the predicted surface using sphere tracing [29]. At each sample point $\boldsymbol{x}_{i}$ and predicted surface distance $s\!=\!f(\boldsymbol{x}_{i})$, we advance by $0.9s$ (chosen empirically to account for our model behaving as an approximate SDF) until hitting the surface (predicted as $s\!\leq\!2\!\times\!10^{-4}$). We only perform sphere tracing where our model behaves as a valid SDF (determined by $\beta(\boldsymbol{x}_{i})\!>\!350$ in our experiments), and fall back to a predefined step size of 1 cm otherwise.

### v2

@@ Abstract
Neural radiance fields provide state-of-the-art view synthesis quality but tend to be slow to render. One reason is that they make use of volume rendering, thus requiring many samples (and model queries) per ray at render time. Although this representation is flexible and easy to optimize, most real-world objects can be modeled more efficiently with surfaces instead of volumes, requiring far fewer samples per ray. This observation has spurred considerable progress in surface representations, such as signed distance functions, but these may struggle to model semi-opaque and thin structures. We propose a method, HybridNeRF, that leverages the strengths of both representations by rendering most objects as surfaces while modeling the (typically) small fraction of challenging regions volumetrically. We evaluate HybridNeRF against the challenging Eyeful Tower dataset [38] along with other commonly used view synthesis datasets. When comparing to state-of-the-art baselines, including recent rasterization-based approaches, we improve error rates by 15–30% while achieving real-time framerates (at least 36 FPS) for virtual-reality resolutions (2K×$\times$2K). Project page: https://haithemturki.com/hybrid-nerf/.
@@ 1 Introduction
Figure 2: Approach. In the first phase of our pipeline (a), we train a VolSDF-like [39] model with distance-adjusted Eikonal loss to model backgrounds without a separate NeRF (Sec. 3.3). We then crucially transition from a uniform surfaceness parameter $\beta$ to position-dependent $\beta(\boldsymbol{x})$ values to model most of the scene as thin surfaces (needing few samples) without degrading quality near fine and semi-opaque structures (b). Since our model behaves as a valid SDF in $>$95% of the scene, we use sphere tracing at render time (c) along with lower-level optimizations (hardware texture interpolation) to query each sample as efficiently as possible.
@@
While surface-based neural fields are convenient for rendering, they often struggle to reconstruct scenes with thin structures or view-dependent effects, such as reflections and translucency. This is one reason that surfaces are often transformed into volumetric models for rendering [39]. A crucial transformation parameter is a scalar temperature $\beta$ that is used to convert a $\beta$-scaled signed distance value into a density. Higher temperatures tend to produce an ideal binary occupancy field that can improve rendering speed but can struggle for challenging regions as explained above. Lower temperatures allow the final occupancy field to remain flexible, whereby the $\beta$-scaled SDF essentially acts as a reparameterization of the underlying occupancy field. As such, we refer to $\beta$ as the surfaceness of the underlying scene (see Fig. 1). Prior work treats $\beta$ as a global parameter that is explicitly scheduled or learned via gradient descent [39]. We learn it in a spatially adaptive manner.
@@ 2 Related Work
Some of the earliest NeRF acceleration methods store precomputed non-view dependent model outputs, such as spherical harmonics coefficients, into finite-resolution structures [43, 13, 9, 6, 7]. These outputs are combined with viewing direction to compute the final radiance at render time, bypassing the original model entirely. Although these methods render extremely quickly (some $>$200 FPS [9]), they are limited by the finite capacity of the caching structure and cannot capture fine details at room scale.
@@
Several methods [25, 35, 39] derive density values from the outputs of a signed distance function, which are then rendered volumetrically as in NeRF. These hybrid representations retain NeRF’s ease of optimization while improving surface geometry. Follow-ups [40, 10] bake the resulting surface geometry into a mesh that is further optimized and simplified. Similar to early voxel-baking approaches, these methods render quickly ($>$70 FPS) but are limited by the capacity of the mesh and texture, and thus struggle to model thin structures, transparency, and view-dependent effects. We train a similar SDF representation in our method but continue using the base neural model at render time. Concurrent to our work, Adaptive shells [37] augments NeuS [35] with a spatially-varying kernel similar to our adaptive surfaceness described in Sec. 3.2.
@@ 3 Method
Given a collection of RGB images and camera poses, our goal is to learn a 3D representation that generates novel views at VR resolution (at least 2K×$\times$2K pixels) in real-time (at least 36 FPS), while achieving a high degree of visual fidelity. As we target captures taken under real-world conditions, our representation must be able to account for inconsistencies across training images due to lighting changes and shadows (even in “static" scenes). We build upon NeRF’s raycasting paradigm, which can generate highly photorealistic renderings, and improve upon its efficiency. As the world mostly consists of surfaces, we train a representation that can render surfaces with few samples and without degrading the rest of the scene. We outline our method in Fig. 2 and present our model architecture and the first training stage in Sec. 3.1, which is followed by finetuning of our model to accelerate rendering without compromising quality in Sec. 3.2. We discuss how to model unbounded scenes in Sec. 3.3 and present final render-time optimizations in Sec. 3.4.
@@ 3.1 Representation
NeRF [22] represents a scene as a continuous volumetric radiance field that encodes the scene’s geometry and view-dependent appearance within the weights of an MLP. NeRF renders pixels by sampling positions $\boldsymbol{x}_{i}$ along the corresponding camera ray, querying the MLP to obtain density and color values, $\sigma_{i}\coloneqq\sigma(\boldsymbol{x}_{i})$ and $\mathbf{c}_{i}\coloneqq\boldsymbol{c}(\boldsymbol{x}_{i},\boldsymbol{d}_{r})$, respectively (with $\boldsymbol{d}_{r}$ as the ray direction). The density values $\sigma_{i}$ are converted into opacity values $\alpha_{i}\coloneqq 1-\exp(-\sigma_{i}\delta_{i})$, where $\delta_{i}$ is the distance between samples. The final ray color $\hat{\boldsymbol{c}}_{r}\coloneqq\sum_{i=0}^{N-1}\boldsymbol{c}_{i}w_{i}$ is obtained as the combination of the color samples $\boldsymbol{c}_{i}$ with weights $w_{i}\coloneqq\exp(-\!\sum_{j=0}^{i-1}\!\sigma_{j}\delta_{j})\alpha_{i}$. The training process optimizes the model by sampling batches of image pixels and minimizing the L2 reconstruction loss. We refer to Mildenhall et al. [22] for details.
@@
Figure 3: Surfaces. Since NeRF directly predicts density, it often ‘cheats’ by modeling specular surfaces, such as floors, as semi-transparent volumes that require many samples per ray (heatmaps shown on the right, with brighter values corresponding to more samples). Methods that derive density from signed distances, such as ours, improve surface geometry and appearance while using fewer samples per ray.
@@
The original NeRF representation has the flexibility of representing semi-transparent surfaces, for the density field is not forced to saturate. However, the model often abuses this property by generating semi-transparent volumes to mimic reflections and other view-dependent effects (Fig. 3). This hampers our goal of minimizing the samples per ray needed for rendering.
To address this problem, surface–volume representations [25, 39, 35] learn well-defined surfaces by interpreting MLP outputs $f(\boldsymbol{x})$ as a signed distance field (SDF) to represent scene surfaces as the zero-level set of the function $f$.
@@
where $\beta(\boldsymbol{x})>0$ determines the surfaceness of point $\boldsymbol{x}$, i.e. how concentrated the density should be around the zero-level set of $f$, and $\Psi$ is the CDF of a standard Laplace distribution:
@@
In prior works, the surfaceness $\beta(\boldsymbol{x})$ is independent of position $\boldsymbol{x}$. We instead consider a surfaceness field implemented as a $512^{3}$ grid of values queried via nearest-neighbor interpolation. We first constrain the surfaceness parameters to be globally uniform, and allow them to diverge spatially during the finetuning stage (Sec. 3.2).
@@
We use dense multi-resolution 3D feature grids in combination with multi-resolution triplanes [4, 8] to featurize 3D sample locations. We predict color $\boldsymbol{c}$ and signed distance $f$ with separate grids, each followed by an MLP, and use a small proposal network similar to that used by Nerfacto [33] to improve sampling efficiency. For a given 3D point, we fetch $K\!=\!4$ features per level from (1) the 3D feature grids at 3 resolution levels (1283, 2563 and 5123) via trilinear interpolation, and (2) from triplanes at 7 levels (from 1282 to 8,1922) via bilinear interpolation. We sum the features across levels (instead of concatenation [23, 8]), and concatenate the summed features from the 3D grid to those from the 3 triplanes to obtain a $164K\!=\!$ 16-dimensional MLP input. We encode viewing direction through spherical harmonics (up to the 4th degree) as an auxiliary input to the color MLP. As our feature grid is multi-resolution, we handle aliasing as in VR-NeRF [38]. See Appendix B and Appendix C for more details.
@@ 3.2 Finetuning
The first stage of our pipeline uses a global surfaceness value $\beta(\boldsymbol{x})=\bar{\beta}$ for all $\boldsymbol{x}$, as in existing approaches [35, 39]. As $\bar{\beta}$ increases, the density $\sigma_{\text{SDF}}$ in free-space areas converges to zero (Eq. 3), reducing the required number of samples per ray. However, uniformly increasing this scene-wide parameter degrades the rendering quality near fine-grained and transparent structures (see Fig. 4).
@@
Figure 5: Spatially adaptive surfaceness. We make $\beta(\boldsymbol{x})$ spatially adaptive by means of a $512^{3}$ voxel grid that we increase during the finetuning stage. We track Eikonal loss as we increase surfaceness as it is highest near object boundaries and semi-transparent surfaces (top-right, brighter = higher loss) that degrade when surfaceness is too high (Fig. 4). We stop increasing surfaceness in regions that cross a given threshold.
We overcome this limitation by making $\beta(\boldsymbol{x})$ spatially adaptive via a $512^{3}$ voxel grid. One possible approach is to directly optimize $\beta(\boldsymbol{x})$ via gradient descent, but we find that this overly relaxes the constraint on SDF correctness such that $f(\boldsymbol{x})$ predicts arbitrary density values as in the original NeRF. We instead rely on the Eikonal loss as a natural indicator of where the model cannot accurately reconstruct the scene via an SDF (and where we should therefore use a “softer” formulation). We collect per-sample triplets $(\boldsymbol{x},\eta,w)$ rendered during the finetuning process, accumulate them over multiple training iterations (5,000), and partition them across the voxels of the surfaceness grid. Let $\Lambda_{\boldsymbol{v}}$ be the subset associated with voxel $\boldsymbol{v}$ corresponding to $\beta_{\boldsymbol{v}}$. We increase $\beta_{\boldsymbol{v}}$ by a fixed increment (100) if:
@@
where $\bar{\gamma}\coloneqq 0.25$ is a predefined threshold. Fig. 5 illustrates our approach.
@@
We find it important to use a large 256 channel-wide MLP to represent the signed distance $f$ during the first training phase in order to learn accurate scene geometry. However, we later distill $f$ into a smaller 16-wide network ($f_{\text{small}})$. We do so by sampling random rays from our training set for 5,000 iterations and minimizing the difference between $f(\boldsymbol{x}_{i})$ and $f_{\text{small}}(\boldsymbol{x}_{i})$ at every sampled point:
@@ 3.3 Backgrounds
Many scenes we wish to reconstruct contain complex backgrounds that surface–volume methods struggle to replicate [39, 35, 19]. BakedSDF [40] defines a contraction space [2] in which the Eikonal loss of Eq. 1 is applied. However, we found this to negatively impact foreground quality. Other approaches use separate NeRF background models [45], which effectively doubles inference and memory costs, and makes them ill-suited for real-time rendering.
@@
We discuss how to make a single MLP behave as an approximate SDF in the foreground and a volumetric model in the background. Both types of NeRF derive density $\sigma$ by applying a non-linearity to the output of an MLP. Our insight is that although the original NeRF uses ReLU, any non-linear mapping to $\mathbb{R}^{+}$ may be used in practice, including our scaled CDF $\Psi$ ($\beta$ omitted without loss of generality). Since $\Psi$ is invertible (as it is a CDF), $\sigma(\boldsymbol{x})$ and $\Psi(f(\boldsymbol{x}))$ are functionally equivalent as there exists an $f$ such that $\Psi(f(\boldsymbol{x}))$ = $\sigma(\boldsymbol{x})$ for any given point $\boldsymbol{x}$. Put otherwise, it is the Eikonal regularization that causes the divergence in behavior between both methods — in its absence, an “SDF" MLP is free to behave exactly as the density MLP in the original NeRF!
@@
We use a distance-adjusted Eikonal loss during training by using per-sample loss weights $\eta_{i}=\frac{1}{d_{i}^{2}}$ (where $d_{i}$ is the metric distance along the ray of sample $\boldsymbol{x}_{i}$) instead of commonly-used uniform weights ($\eta_{i}=1$) to downweight the loss applied to far-field regions. Intuitively, this encourages our method to behave as a valid SDF in the foreground (with well-defined surfaces) and more like NeRF in the background (to enable accurate reconstruction) without the need for separate foreground and background models. Fig. 6 and Tab. 7 illustrate the different approaches.
@@ 3.4 Real-Time Rendering
Volumetric methods that use occupancy grids [e.g. 23, 29] sample within occupied voxels using a given step size. This hyperparameter must be carefully tuned to strike the proper balance between quality (not skipping thin surfaces) and performance (not excessively sampling empty space). Modeling an SDF allows us to sample more efficiently by advancing toward the predicted surface using sphere tracing [31]. At each sample point $\boldsymbol{x}_{i}$ and predicted surface distance $s\!=\!f(\boldsymbol{x}_{i})$, we advance by $0.9s$ (chosen empirically to account for our model behaving as an approximate SDF) until hitting the surface (predicted as $s\!\leq\!2\!\times\!10^{-4}$). We only perform sphere tracing where our model behaves as a valid SDF (determined by $\beta(\boldsymbol{x}_{i})\!>\!350$ in our experiments), and fall back to a predefined step size of 1 cm otherwise.